# EF-3DGS: Event-Aided Free-Trajectory 3D Gaussian Splatting

### v1

@@ V-C1 Metrics
[…]
EvDeblurNeRF $\times$ E+F 20.24 0.5762 19.35 0.5649 20.47 0.5757 22.36 0.7087 18.23 0.5043 16.41 0.4624 17.82 0.4959 19.44 0.6448
ENeRF $\times$ E+F 21.89 0.5933 21.94 0.5918 22.01 0.5918 23.91 0.7266 20.09 0.5261 18.99 0.4903 19.61 0.5173 20.00 0.6494
[…]
TABLE IV: Pose accuracy on RealEv-DAVIS. The unit of $\textrm{RPE}_{r}$ is in degrees, ATE is in the ground truth scale and $\textrm{RPE}_{t}$ is scaled by 100. Those methods which require precomputed poses are excluded.
Methods Input SLOW FAST
hall building corner outdoor hall building corner outdoor
$\textrm{RPE}_{t}$$\downarrow$ $\textrm{RPE}_{r}$$\downarrow$ $\textrm{RPE}_{t}$$\downarrow$ $\textrm{RPE}_{r}$$\downarrow$ $\textrm{RPE}_{t}$$\downarrow$ $\textrm{RPE}_{r}$$\downarrow$ $\textrm{RPE}_{t}$$\downarrow$ $\textrm{RPE}_{r}$$\downarrow$ $\textrm{RPE}_{t}$$\downarrow$ $\textrm{RPE}_{r}$$\downarrow$ $\textrm{RPE}_{t}$$\downarrow$ $\textrm{RPE}_{r}$$\downarrow$ $\textrm{RPE}_{t}$$\downarrow$ $\textrm{RPE}_{r}$$\downarrow$ $\textrm{RPE}_{t}$$\downarrow$ $\textrm{RPE}_{r}$$\downarrow$
LocalRF F 4.4017 2.5067 4.6186 1.9409 1.9296 2.5183 3.4307 1.3005 5.6061 3.2265 7.0611 2.8722 3.4767 3.8950 4.7229 1.8595
CF-3DGS F 2.7106 0.6035 3.0374 3.2761 1.9437 1.4954 2.2747 0.8053 3.5025 0.7639 4.6758 4.5256 3.4950 2.2855 3.0551 1.1198
Ev-Baseline E+F 0.7063 0.4386 3.4225 1.1717 2.1283 1.3361 0.8647 0.3455 0.8898 0.7734 5.0707 1.6997 3.6785 2.1552 1.1645 0.5001
EF-3DGS(Ours) E+F 0.5262 0.3251 2.8008 0.4633 1.8386 1.6168 0.4726 0.3699 0.5424 0.3464 3.5848 0.5589 2.5396 2.2157 0.5218 0.4462
V-C2 Baselines
For a fair comparison, we focus on two categories of methods: those solely based on frames and those leveraging both events and frames. For frame-based approaches, we selected methods specifically addressing free-trajectory scenarios, such as LocalRF and F2-NeRF, although the latter requires pre-estimated poses. We also include pose-free methods like Nope-NeRF and the current state-of-the-art CF-3DGS. To address the out-of-memory (OOM) issues CF-3DGS faces with long video sequences, we implement a slight modification, segmenting long videos into subsequences for individual reconstruction. For event-based methods, consistent with EvDeblurNeRF [47], we consider approaches that fuse events and frames. This category includes ENeRF [43] and the SOTA deblurring method EvDeblurNeRF [47]. These hybrid approaches leverage the complementary strengths of both data modalities. To the best of our knowledge, no existing method has exploit event cameras to address free-trajectory problem. To facilitate a fair comparison, we construct an event-based baseline method built upon CF-3DGS [5]. It leverages the high temporal resolution property of event cameras by utilizing an event-based frame interpolation network Time Lens [63] to interpolate the original video. Then, the temporally upsampled video frames are subsequently fed into CF-3DGS.
V-D Experimental Results
To evaluate the results on synthesized novel views, we select every ten frames as a test image following LocalRF [9]. Since the camera poses are unknown in our setting, we need to estimate the poses of test views. As in NeRFmm [40] and iNeRF [64], we freeze the 3DGS model, initialize the test poses with the poses of the nearest training frames, and optimize the test poses by minimizing the photometric error between rendered images and test views.
V-D1 Results on Tanks and Temples
Quantitatively, the experimental results in Table I and Table II demonstrate the superiority of event-based methods (EF-3DGS and Ev-Baseline) over frame-based baselines in both pose estimation accuracy and rendering quality. This highlights the potential of event cameras for free-trajectory scene reconstruction. Furthermore, our method surpasses Ev-Baseline across all evaluation metrics under varying frame rate settings, indicating that our proposed method can better effectively integrate property of the high temporal resolution and motion information encoded in the event streams into the scene reconstruction process.
Qualitatively, as shown in Fig. 4, our method generates more realistic and detailed renderings compared to the baseline. The synthesized views of our method exhibit sharper edges and better preservation of finer textures. In contrast, the baseline methods suffer from various artifacts, such as blurriness and texture inconsistencies. Moreover, the visualization of estimated camera trajectories in Fig. 5 indicates that our estimated trajectory is better aligned with the ground truth than other methods, and the $\textrm{RPE}_{r}$ is significantly smaller.
Figure 6: Qualitative comparison for novel view synthesis on RealEv-DAVIS. Our approach produces more realistic rendering results with fine-grained details. Better viewed when zoomed in.
Figure 7: Pose Estimation Comparison on RealEv-DAVIS. We visualise the trajectory (3D plot) and relative rotation errors $\textrm{RPE}_{r}$ (color bar) of each method. We clip and normalize the $\textrm{RPE}_{r}$ by a quarter of the max $\textrm{RPE}_{r}$ across all results of each scene.
V-D2 Results on RealEv-DAVIS
In Table III and Table IV, we report results on the self-collected dataset, RealEv-DAVIS. We select the top-performing methods from the previous evaluation. EF-3DGS demonstrates robust generalization to real-world scenarios, despite challenges such as sensor noise absent in simulated environments. Notably, the performance gap widens in FAST scenarios, with PSNR improvements of nearly 2 dB over baselines, emphasizing the method’s efficacy in handling high-speed motion in real-world settings. Qualitative results in Fig. 6 and Fig. 7 showcase EF-3DGS’s superior preservation of fine-grained details and the method’s robustness to real-world trajectory.
V-D3 Performance under Varying Camera Speeds
As evidenced in Table I to Table IV, the performance of all methods generally declines as the frame rate decreases. However, our method exhibits remarkable resilience to the reduction in frame rate, maintaining its superiority over the other methods across all metrics. Regarding novel view synthesis quality, the performance gap between the proposed method and the other methods consistently widens as the frame rate decreases. For example, the rendering quality difference between our method and CF-3DGS [5], measured in PSNR, increases from 0.61dB at 6FPS to 3.43dB at 1FPS. This observation demonstrates the robustness of EF-3DGS in handling challenging high-speed scenarios, where the advantages of leveraging event data become more pronounced.
V-D4 Robustness to Pose Disturbance
To further validate the performance of different methods under inaccurate pose initialization, a common challenge in practical scenarios, we introduce varying degrees of perturbations to the initial camera poses estimated by COLMAP. Specifically, following BARF [42], we parametrize the camera poses $\mathbf{p}$ with the $\mathfrak{se}(3)$ Lie algebra. For each scene, we synthetically perturb the camera poses with additive noise $\delta\mathbf{p}\sim\mathcal{N}(\mathbf{0},n\mathbf{\textrm{I}})$, where n is the noise level. Then, each method is initialized with the noised poses, after which the optimization is performed. The results are illustrated in Fig. 11. Notably, ENeRF [43], which lacks the capability to optimize camera poses, exhibits a drastic performance degradation as the magnitude of pose disturbances increases. This observation validates the critical importance of optimizing camera poses alongside the scene reconstruction. Furthermore, Our proposed framework exhibits superior tolerance across the entire range of pose perturbations compared to existing methods. Even in the presence of significant perturbation, the degradation in EF-3DGS’s novel view synthesis fidelity and trajectory estimation performance is substantially lower than that of the other methods.
Figure 8: Qualitative Results of the effectiveness of $\mathcal{L}_{EGM}$.
TABLE V: Effect of each component in EF-3DGS. Evaluation is conducted on 1FPS scenes of Tanks and Temples. The best results are highlighted in bold.
@@ V-D4 Robustness to Pose Disturbance
Figure 9: Qualitative Results of the effectiveness of $\mathcal{L}_{LEGM}$.
Figure 10: Qualitative Results of the effectiveness of $\mathcal{L}_{PBA}$.
@@
TABLE VI: Effect of the number of event subintervals N.
[…]
$N=2$ 23.56 30+ 1.3h
$N=3$ 24.43 30+ 3h
$N=6$ 24.81 30+ 7h
[…]
CF-3DGS 22.08 30+ 3h
TABLE VII: Comparison of training time vs PSNR for various methods
V-E Ablation Studies
To further investigate the effectiveness of each key component in our proposed EF-3DGS framework, we conduct ablation experiments under the 1 FPS setting on the Tanks and Temples dataset. The results are presented in Table V.
V-E1 Effect of $\mathcal{L}_{EGM}$
Analysis of Table V demonstrates that removing all components substantially degrades both rendering quality and pose estimation accuracy. This finding underscores the significant challenges faced by the original CF-3DGS in high-speed scenarios. Fig. 8 further illustrates this issue, revealing that while CF-3DGS can produce high-quality renderings at training views, it struggles with adjacent novel views. Moreover, the rendered depth maps expose considerable inaccuracies in CF-3DGS’s scene geometry reconstruction. The incorporation of $\mathbf{\mathcal{L}_{EGM}}$ yields marked improvements in both rendering quality and pose estimation precision, as evidenced by Table V and Fig. 8. This experimental outcome suggests that $\mathbf{\mathcal{L}_{EGM}}$ effectively harnesses the continuous brightness change data captured by event cameras, thereby compensating for the inter-frame loss of scene brightness information. Consequently, the 3DGS training optimization achieves multi-view geometric consistency. Notably, among all components, $\mathbf{\mathcal{L}_{EGM}}$ demonstrates the most profound impact on overall method performance, establishing its role as a fundamental component of our approach.
V-E2 Effect of $\mathcal{L}_{LEGM}$
As demonstrated in Table V, $\mathbf{\mathcal{L}_{LEGM}}$ further enhances both rendering quality and pose estimation accuracy when applied in conjunction with $\mathbf{\mathcal{L}_{EGM}}$. Fig. 9 provides a visual illustration of $\mathbf{\mathcal{L}_{LEGM}}$’s efficacy. Ideally, with accurate depth and pose estimation, the Image of Warped Events (IWE) should resemble the reference image. However, in the absence of $\mathbf{\mathcal{L}_{LEGM}}$, the generated IWE exhibits blurring effects, indicative of inaccurate pose estimation. More critically, as highlighted in the boxed area of Fig. 9, the presence of self-similar regions with repetitive textures poses additional challenges for scene reconstruction and pose estimation. Significant errors in scene reconstruction leading to inaccurate depth estimates further compromise the inherent structure of the IWE. The introduction of $\mathbf{\mathcal{L}_{LEGM}}$ allows for the simultaneous reflection of both pose estimation and scene reconstruction errors in the IWE. By optimizing the sharpness of the IWE, $\mathbf{\mathcal{L}_{LEGM}}$ effectively improves both rendering quality and pose estimation precision.
V-E3 Effect of $\mathcal{L}_{PBA}$ and Fixed-GS Strategy
Comparing the last three rows of Table V, we observe that both $\mathbf{\mathcal{L}_{PBA}}$ and Fixed-GS strategies significantly enhance rendering quality. $\mathbf{\mathcal{L}_{PBA}}$, by enforcing multi-view consistency, slightly improves pose estimation accuracy, while the Fixed-GS strategy, applied solely during scene optimization, has no impact on pose estimation. Fig. 10 visualizes the effects of these components. When both are removed, the reconstructed scene loses almost all color information. This is attributed to the substantially higher number of event frames than image frames, causing event-guided optimization to dominate scene reconstruction and suppress color information from image frames. $\mathbf{\mathcal{L}_{PBA}}$ alone, which reprojects neighboring video frames onto event frames, only partially mitigates this imbalance. The Fixed-GS strategy employs a two-stage scene optimization approach for the separation of the color and structure of the scene. However, this approach does not fully leverage the supervisory signals from adjacent frames. The combination of $\mathbf{\mathcal{L}_{PBA}}$ and Fixed-GS strategies complements each other, more effectively addressing the color distortion issue.
V-E4 Computational Efficiency and Impact of Subinterval Number N
To investigate the impact of the number of subintervals N on our method, we conduct ablation studies on the Tanks and Temples dataset under the 2 FPS setting. The results are shown in Table VII and figure on the right. We test the speed on RTX2080ti. Note that in the t-PSNR figure, circular nodes represent pose-free methods, while diamond-shaped nodes indicate methods that require precomputed poses. The results in Table VII reveal that increasing N improves PSNR, indicating that finer temporal discretization enhances reconstruction quality. However, this improvement comes at the cost of extended training time, increasing from 1.3 hours at N=2 to 7 hours at N=6. Notably, our method maintains real-time rendering performance, thanks to the efficient rendering capabilities of 3DGS. Even at N=2, our approach outperforms baselines in both PSNR and rendering speed. The setting of N=3 offers a good balance, achieving higher PSNR than CF-3DGS with comparable training time. However, it’s important to note that the optimal N may vary for different frame rates and motion characteristics.
Figure 12: Study of the different evalution metrics with respect to $\lambda_{cm}$
Figure 13: Study of the different evalution metrics with respect to $\lambda_{grad}$
Figure 14: Study of the different evalution metrics with respect to $\lambda_{PBA}$
V-F Analysis of Loss Coefficients
Different components of our method play crucial roles in overall performance. To comprehensively understand these effects and determine optimal weight settings, we conducted a detailed analysis of the loss weights: $\lambda_{cm}$, $\lambda_{grad}$ and $\lambda_{PBA}$.
V-F1 Contrast Maximization Coefficient $\lambda_{cm}$
As shown in Fig. 14, the contrast maximization coefficient $\lambda_{cm}$ significantly affects both NVS quality and pose estimation. We observe that a value of 0.1 achieves the best overall performance. Lower values (0.01) lead to decreased performance, likely due to insufficient utilization of motion information from events. Higher values (0.5) also result in performance degradation, possibly due to over-reliance on the contrast maximization term at the expense of other constraints. The optimal $\lambda_{cm}$ suggests that while event data is crucial, it should not dominate the reconstruction process. This balance allows our method to leverage the high temporal resolution of events without sacrificing the global supervision provided by frame data.
V-F2 Gradient-based Loss Coefficient $\lambda_{grad}$
As shown in Fig. 14, the gradient-based loss coefficient $\lambda_{grad}$ shows optimal performance at 0.2. Lower values (0.1) slightly decrease performance, while higher values (0.3, 0.5) lead to more significant drops in both NVS quality and pose accuracy. This is likely because $\mathcal{L}_{EGM}$ and 3DGS primarily focus on rendering absolute brightness, and excessive gradient constraints may interfere with this process. Therefore, a moderate $\lambda_{grad}$ value is crucial to balance gradient information with the primary rendering objectives.
V-F3 Photometric Bundle Adjustment Coefficient $\lambda_{PBA}$
As illustrated in Figure 14, our method demonstrates remarkable stability across $\lambda_{PBA}$ values ranging from 0.1 to 0.5. However, Fig. 10 illustrates that while $\mathcal{L}_{PBA}$ has minimal effect on PSNR and ATE metrics, its absence leads to noticeable color distortion artifacts in visual results. This underscores the importance of $\lambda_{PBA}$ in maintaining visual fidelity. Conversely, setting $\lambda_{PBA}$ to 1.0 results in significant performance degradation, indicating that overemphasis on cross-view consistency can be counterproductive.
VI Limitations
Our method combines event streams with conventional video frames, which require the input data to be time-ordered. So it can not handle the unordered image input. In high-speed situations, video frames may suffer from severe motion blur, which can degrade the quality of the visual information. Future research could focus on this challenge. And, our method relies on several key parameters (e.g., $\lambda_{cm}$, $\lambda_{grad}$ and $\lambda_{PBA}$) that may require manual tuning for optimal performance across different speed scenarios. This dependency on scene-specific parameter settings could limit the method’s adaptability to diverse environments. Future work should explore self-adaptive parameter adjustment strategies to enhance the method’s versatility and ease of use across various reconstruction tasks.
VII Conclusions

### v3

@@ 5.4 Experimental Results
LocalRF F 20.83 0.6074 3.60 2.07 17.62 0.5192 5.22 2.96
CF-3DGS F 22.68 0.6287 2.49 1.55 17.59 0.5204 3.68 2.17
EvDeblurNeRF E+F 20.61 0.6064 - - 17.98 0.5269 - -
Event-3DGS (E+F) E+F 23.43 0.6456 - - 20.04 0.5515 - -
EvCF-3DGS E+F 22.89 0.6317 1.78 0.82 19.13 0.5380 2.70 1.28
EF-3DGS(Ours) E+F 23.65 0.6466 1.41 0.69 21.12 0.5620 1.80 0.89
Table 3: Rendering and pose estimation results on RealEv-DAVIS. Complete data and additional metrics are provided in the supplementary material.
Results on RealEv-DAVIS.  Table 3 validates our approach on the real-world RealEv-DAVIS dataset. EF-3DGS outperforms top-performing methods and handles real-world scenes effectively. In FAST scenarios, our method shows nearly 1dB PSNR improvement over the best baselines. This confirms our advantage in high-speed scenarios where frame-based methods struggle. Fig. 4 and Fig. 5 show our method preserves fine details and maintains accurate trajectories even during rapid motion, addressing key limitations of traditional approaches.
Performance under Varying Camera Speeds  As shown in Fig 2 and Fig. 3, while all methods degrade as the frame rate decreases, our approach shows remarkable resilience. The performance gap widens significantly at lower frame rates, with our PSNR advantage over CF-3DGS [9] increasing from 0.61dB at 6FPS to 3.43dB at 1FPS. Notably, our method also consistently outperforms other event-based methods (EvCF-3DGS and Event-3DGS). This confirms not only the value of event data in challenging scenarios but also the superiority of our integration approach.
@@
Table 4: Effect of each component in EF-3DGS. The best results are highlighted in bold.
5.5 Ablation Studies
Effect of Each Component  Table 4 presents a comprehensive ablation study of our key components under the challenging 1FPS setting on Tanks and Temples. $\mathcal{L}_{EGM}$ serves as the foundation of our approach, providing substantial improvements in both rendering quality (+1.63dB PSNR) and pose accuracy by enabling rich supervision between discrete frames. Building upon this, $\mathcal{L}_{LEGM}$ extracts motion information from events and constrains 3DGS in the gradient domain, significantly improving pose estimation while modestly enhancing rendering quality. $\mathcal{L}_{PBA}$, though designed to address color inconsistency issues, not only improves rendering quality but also enhances pose estimation accuracy by establishing geometric and photometric consistency across views. The Fixed-GS training strategy, while having no impact on pose optimization, significantly improves rendering quality by effectively separating structure and color optimization. We provide more intuitive ablation visualizations in the supplementary material.
@@ 5.5 Ablation Studies
Robustness to Pose Disturbance  To validate the robustness of different methods under inaccurate pose initialization, a common challenge in practical scenarios, we introduce varying degrees of perturbations to the initial camera poses estimated by COLMAP. Specifically, following BARF [23], we parametrize the camera poses $\mathbf{p}$ with the $\mathfrak{se}(3)$ Lie algebra. For each scene, we synthetically perturb the camera poses with additive noise $\delta\mathbf{p}\sim\mathcal{N}(\mathbf{0},n\mathbf{\textrm{I}})$, where n is the noise level. Then, each method is initialized with the noised poses, after which the optimization is performed. The results are illustrated in Fig. 6. Notably, Event-3DGS [16], which lacks the capability to optimize camera poses, exhibits a drastic performance degradation as the magnitude of pose disturbances increases. This observation validates the critical importance of joint pose-scene optimization. Furthermore, Our proposed framework demonstrates superior tolerance across all perturbation levels. Even under significant noise, our method experiences substantially less degradation in both rendering quality and trajectory accuracy.
6 Conclusions